# A Nonconforming Virtual Element Method for Advection-Diffusion-Reaction Problems with CIP Stabilization

### v1

@@ 1 Introduction
{comment}
[OLD:
  • extension to nonconforming of your previous paper
  • advantages of nonconforming
  • use of the standard variational formulation (no skew-symmetrization)
  • symmetric Nitsche: natural for diffusion problems; numerical advantages?
  • why not also in 3D?
@@ 2.1 Continuous Problem
Given a domain $\Omega\subset\mathbb{R}^{2}$ with a polygonal boundary $\Gamma$, we consider the steady advection-diffusion-reaction equation with non-homogeneous Dirichlet boundary condition
@@
For simplicity, we assume that the diffusion and the reaction coefficient $\varepsilon$ and $\sigma$, respectively, are two positive constants, while the advection coefficient $\boldsymbol{\beta}\in[W^{1,\infty}(\Omega)]^{2}$ is such that $\text{div}(\boldsymbol{\beta})=0$. Finally, $f\in L^{2}(\Omega)$ and $g\in H^{\frac{1}{2}}(\Gamma)$. We introduce the Sobolev space
@@
where $\boldsymbol{n}$ is the outward unit normal vector to the boundary. Integrating by parts the first term in (1), we derive a weak formulation for problem (1):
@@
$\|\boldsymbol{\beta}\|_{[L^{\infty}(\Omega)]^{2}}=1\,.$ (7)
@@ 2.2 Notation and Polynomial Approximation Results
We consider a sequence $\{\Omega_{h}\}_{h}$ of tessellations of the domain $\Omega$ into non-overlapping polygons. From now on, let $E$ denote an (open) element of the mesh $\Omega_{h}$, and $e$ an edge of the polygon $E$. Moreover, let $|E|$, $h_{E}$, and $\boldsymbol{n}^{E}$ denote the area, the diameter, and the unit outward normal of the polygon $E$, respectively. We also write $\boldsymbol{n}^{e}$ for the restriction of $\boldsymbol{n}^{E}$ to the edge $e\subset\partial E$. We define the mesh size as $h\coloneqq\sup_{E\in\Omega_{h}}h_{E}$.
The set of the edges of a tessellation $\Omega_{h}$ is denoted with $\mathcal{E}_{h}$. This set is split in internal edges and boundary edges $\mathcal{E}_{h}=\mathcal{E}_{h}^{0}\cup\mathcal{E}_{h}^{\partial}$. For any scalar-valued function $v$, whose restriction $v^{E}$ to each element $E\in\Omega_{h}$ is sufficiently smooth, we define the (vector-valued) jump $[\![v]\!]$ and the (scalar-valued) average $\{v\}$ on each edge $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$ as
@@
Similarly, for any piecewise smooth vector-valued function $\boldsymbol{v}$, we define the (scalar-valued) jump $[\![\boldsymbol{v}]\!]$ on each edge $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$ as
@@
  • any edge $e$ of $E$ has length $\geq\,\varrho\,h_{E}$;
  • the mesh is quasi-uniform, i.e., any polygon has diameter $h_{E}\geq\varrho h$.
We introduce some basic spaces that will be useful later on. Given two positive integers $n$ and $m$, and $p\in[0,+\infty]$, for any $E\in\Omega_{h}$ we define
@@
    with obvious extensions for functions defined on an edge $\Pi_{n}^{0,e}\colon L^{2}(e)\to\mathbb{P}_{n}(e)$, and for vector-valued functions $\boldsymbol{\Pi}^{0,E}_{n}\colon[L^{2}(E)]^{2}\to[\mathbb{P}_{n}(E)]^{2}$ and $\boldsymbol{\Pi}^{0,e}_{n}\colon[L^{2}(e)]^{2}\to[\mathbb{P}_{n}(e)]^{2}$;
@@
$\|\varphi-\Pi^{0,E}_{n}\varphi\|_{W^{m}_{p}(E)}\lesssim h_{E}^{s-m}|\varphi|_{W^{s}_{p}(E)}\qquad$ $s,m\in\mathbb{N}$, $m\leq s\leq n+1$, $p=1,\dots,\infty$,
@@
$\|\nabla\varphi-\boldsymbol{\Pi}^{0,E}_{n}\nabla\varphi\|_{m,E}\lesssim h_{E}^{s-1-m}|\varphi|_{s,E}\qquad$ $s,m\in\mathbb{N}$, $m+1\leq s\leq n+1$, $s\geq 1$.
We also have a counterpart of this result for projections defined on an edge.
@@
Let $e\in\mathcal{E}_{h}^{0}$ be an internal edge and let $E,E^{\prime}\in\Omega_{h}$ be such that $e\subset\partial E\cap\partial E^{\prime}$. Moreover, let $\Pi^{0,e}_{n}$ denote the $L^{2}(e)$-orthogonal projection onto $\mathbb{P}_{n}(e)$. Then, for every sufficiently smooth function $\varphi$ defined on $E\cup E^{\prime}$, we have that
@@ 2.3 Virtual Element Spaces
Given an element $E\in\Omega_{h}$ with $n_{E}$ edges, and a positive integer $k$, we define the space
@@
  • $\mathcal{E}^{E,k}$: the moments up to the order $k-1$ on each edge $e\subset\partial E$:
    $\mu_{E,e}^{l}(v_{h})\coloneqq\dfrac{1}{|e|}\int_{e}v_{h}\left(\dfrac{s-s_{e}}{h_{e}}\right)^{l}{\rm d}s\,,\quad l=0,\ldots,k-1,$ (11)
    where $s$ denotes a local coordinate on $e$, and $s_{e}$ is the coordinate value corresponding to the edge midpoint.
@@
    where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})^{T}$ is a multi-index.
In particular, for $k=1$, the basis functions for the space $V_{h}^{k,nc}(E)$ are associated only with the edge DoFs; for $k\geq 2$, there are also the interior DoFs. In Figure 1, we show the DoFs on a penthagon for $k=1,2,3$.
@@
$N_{E}=k\,n_{E}+\dfrac{(k-1)(k-2)}{2}\,.$
@@
$\int_{e}[\![v_{h}]\!]\cdot\boldsymbol{n}^{e}\,q=0\quad\forall q\in\mathbb{P}_{k-1}(e)\quad\forall e\in{\mathcal{E}}_{h}^{0}\bigr{\}}\,.$
We remark that we do not impose full continuity across the element boundaries. On each interior edge, we only require that the moments up to order $k-1$ are preserved. Therefore, $V_{h}^{k,nc}(\Omega_{h})\not\subset H^{1}(\Omega)$. The global DoFs are as follows:
@@
    $\mu_{e}^{l}(v_{h})\coloneqq\dfrac{1}{|e|}\int_{e}v_{h}\left(\dfrac{s-s_{e}}{h_{e}}\right)^{l}{\rm d}s\,,\quad l=0,\ldots,k-1;$
@@
    where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})^{T}$ is a multi-index.
@@
$N=k\,|\mathcal{E}|+\dfrac{(k-1)(k-2)}{2}\,|\Omega_{h}|\,,$
where $|\mathcal{E}|$ and $|\Omega_{h}|$ are the number of edges and of elements, respectively. Letting $v_{\mathcal{I}}$ be the DoFs-interpolant of a sufficiently smooth function $v$ into $V_{h}^{k,nc}(\Omega_{h})$, we recall the following interpolation error estimate, see [1, 17].
@@
Under assumption (A1), for any $v\in V\cap H^{s+1}(\Omega_{h})$, there exists $v_{\mathcal{I}}\in V_{h}(\Omega_{h})$, such that for all $E\in\Omega_{h}$,
@@ 2.4 Virtual Element Forms and the Discrete Problem
The local bilinear form $a^{E}(\cdot,\cdot)$ is discretized by replacing the gradient with its projection into the space $[\mathbb{P}_{k-1}(E)]^{2}$. We define the bilinear form $a_{h}^{E}(\cdot,\cdot)\colon V_{h}^{k,nc}(E)\times V_{h}^{k,nc}(E)\to\mathbb{R}$ as follows:
$\begin{split}a_{h}^{E}(u_{h},v_{h})&:=\int_{E}{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{h}\cdot{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\,{\rm d}E+\mathcal{S}^{E}\bigl{(}(I-{\Pi^{\nabla,E}_{k}})u_{h},\,(I-{\Pi^{\nabla,E}_{k}})v_{h}\bigr{)}\,{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0},}\end{split}$ (15)
@@
The reasons behind this bilinear form will become clear in Lemma 3.3 below and are related to the nonconformity of the method. Following [16, 15], we also introduce a VEM version of the local CIP-stabilization form, in order to ensure tability in the advection-dominated regime. It is defined as
@@
where $\kappa_{e}$ and $\kappa_{E}$ are positive constants to be chosen. We stress that, whenever $\boldsymbol{\beta}=0$ inside an element $E$, the corresponding parameters $\gamma_{e}$ and $\gamma_{E}$ are zero, and hence $J_{h}^{E}(u_{h},v_{h})=0$. In order to complete the definition of our method, we need to impose the boundary conditions. We do this using a Nitsche-type technique, which locally consists of adding the form
@@
see [26, 22] for more details. Adding these terms together, we define the complete local bilinear form $\mathcal{A}_{\rm cip}^{E}\colon V_{h}^{k,nc}(E)\times V_{h}^{k,nc}(E)\to\mathbb{R}$ as
$\mathcal{A}_{\rm cip}^{E}(u_{h},v_{h})=\varepsilon a_{h}^{E}(u_{h},v_{h})+b^{E}_{h}(u_{h},v_{h})+\sigma c_{h}^{E}(u_{h},v_{h})+d^{E}_{h}(u_{h},v_{h})+\mathcal{N}_{h}^{E}(u_{h},v_{h})+J_{h}^{E}(u_{h},v_{h})\,,$ (23)
and, summing over all the mesh elements, we obtain the global versions of the bilinear forms
@@ 3 Theoretical Analysis
Due to the nonconformity of the virtual element space and the projections entering formulation (26), the solution $u$ of the continuous problem (2) does not solve the discrete problem (26). However, if $u\in H^{2}(\Omega)\cap H_{g}^{1}(\Omega)$, it solves the following problem, which is strictly connected to (26):
@@ 3.1 Preliminary results
Since $\Delta v_{h}$ is polynomial we have (see for instance [6])
$\|\,\Delta v_{h}\|_{0,E}\lesssim h_{E}^{-1}|v_{h}|_{1,E}.$ (33)
Therefore, for the first term on the right-hand side of (38), we get
$-\int_{E}v_{h}\,\Delta v_{h}\,{\rm d}E\leq\|v_{h}\|_{0,E}\|\,\Delta v_{h}\|_{0,E}\lesssim h^{-1}_{E}\,\|v_{h}\|_{0,E}|v_{h}|_{1,E}\,.$ (34)
$\leq C_{t}\|v_{h}\|_{0,E}^{\frac{1}{2}}\,\left(h^{-1}_{E}\|v_{h}\|_{0,E}+|v_{h}|_{1,E}\right)^{\frac{1}{2}}\,\left\|\nabla v_{h}\cdot\boldsymbol{n}^{E}\right\|_{0,\partial E}\,,$
for a positive constant $C_{t}>0$. Since $\nabla v_{h}\cdot\boldsymbol{n}^{E}$ is polynomial on the sides of $\partial E$, a 1D inverse estimate and a trace inequality give
$\left\|\nabla v_{h}\cdot\boldsymbol{n}^{E}\right\|_{0,\partial E}\lesssim h_{E}^{-\frac{1}{2}}\left\|\nabla v_{h}\cdot\boldsymbol{n}^{E}\right\|_{-\frac{1}{2},\partial E}\lesssim h_{E}^{-\frac{1}{2}}|v_{h}|_{1,E}.$ (36)
Hence, we obtain
$\int_{\partial E}v_{h}\,\nabla v_{h}\cdot\boldsymbol{n}^{E}\,{\rm d}s$ $\lesssim C_{t}h^{-\frac{1}{2}}_{E}\|v_{h}\|_{0,E}^{\frac{1}{2}}\,\left(h^{-1}_{E}\|v_{h}\|_{0,E}+|v_{h}|_{1,E}\right)^{\frac{1}{2}}|v_{h}|_{1,E}$ (37)
$\lesssim\left(\alpha+\frac{C_{t}^{2}}{\alpha}\right)h^{-1}_{E}\|v_{h}\|_{0,E}|v_{h}|_{1,E}+\alpha|v_{h}|^{2}_{1,E}\,,$
for any $\alpha>0$. From (38), (40), and (43), we get
$|v_{h}|_{1,E}\lesssim\left(1+\alpha+\frac{C_{t}^{2}}{\alpha}\right)h^{-1}_{E}\|v_{h}\|_{0,E}+\alpha|v_{h}|_{1,E}\,.$
Choosing $\alpha$ sufficiently small, we conclude
$|v_{h}|_{1,E}\lesssim h^{-1}_{E}\|v_{h}\|_{0,E}\,.$
An integration by parts gives
Since $\Delta v_{h}$ is polynomial we have (see for instance [6])
@@
$\pi v=\sum_{e\in\mathcal{E}_{h}}\sum_{l=0}^{k-1}\mu_{e}^{l}(v)\varphi_{e}^{l}+\sum_{E\in\Omega_{h}}\sum_{|\boldsymbol{\alpha}|\leq k-2}\mu^{\boldsymbol{\alpha}}_{E}(v)\varphi_{E}^{\boldsymbol{\alpha}}\,,$ (50)
where $\{\varphi_{e}^{l}\}$ is the set of basis functions associated to the DoFs at the mesh skeleton $\mathcal{E}^{k}$, and $\{\varphi_{E}^{\boldsymbol{\alpha}}\}$ is the set of basis functions associated to the interior DoFs. For an interior edge $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$, the coefficient $\mu_{e}^{l}(v)$ is defined as
$\mu_{e}^{l}(v)\coloneqq\dfrac{1}{2}\left(\mu_{E,e}^{l}(v)+\mu_{E^{\prime},e}^{l}(v)\right),$ (51)

### v2

@@ 2.1 Continuous Problem
Given a polytopal domain $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, with boundary $\Gamma$, we consider the steady advection-diffusion-reaction equation with non-homogeneous Dirichlet boundary condition
@@
For simplicity, we assume that the diffusion and the reaction coefficient $\varepsilon$ and $\sigma$, respectively, are two positive constants, while the advection coefficient $\boldsymbol{\beta}\in[W^{1,\infty}(\Omega)]^{d}$ is such that $\text{div}(\boldsymbol{\beta})=0$. Finally, $f\in L^{2}(\Omega)$ and $g\in H^{\frac{1}{2}}(\Gamma)$. We introduce the Sobolev space
@@
where $\boldsymbol{n}$ is the outward unit normal vector to the boundary $\Gamma$. Integrating by parts the first term in (1), we derive a weak formulation for problem (1):
@@
$\|\boldsymbol{\beta}\|_{[L^{\infty}(\Omega)]^{d}}=1\,.$ (7)
@@ 2.2 Notation and Polynomial Approximation Results
We consider a sequence $\{\Omega_{h}\}_{h}$ of tessellations of the domain $\Omega$ into non-overlapping polytopes. From now on, let $E$ denote an (open) element of the mesh $\Omega_{h}$, and $e$ facet (edge if $d=2$, face if $d=3$) of the element $E$. Moreover, let $|E|$, $h_{E}$, and $\boldsymbol{n}^{E}$ denote the volume, the diameter, and the unit outward normal of the element $E$, respectively. We also write $\boldsymbol{n}^{e}$ for the restriction of $\boldsymbol{n}^{E}$ to the facet $e\subset\partial E$. We define the mesh size as $h\coloneqq\sup_{E\in\Omega_{h}}h_{E}$.
The set of the facets of a tessellation $\Omega_{h}$ is denoted with $\mathcal{E}_{h}$. This set is split in internal facets and boundary facets $\mathcal{E}_{h}=\mathcal{E}_{h}^{0}\cup\mathcal{E}_{h}^{\partial}$. For any scalar-valued function $v$, whose restriction $v^{E}$ to each element $E\in\Omega_{h}$ is sufficiently smooth, we define the (vector-valued) jump $[\![v]\!]$ and the (scalar-valued) average $\{v\}$ on each facet $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$ as
@@
Similarly, for any piecewise smooth vector-valued function $\boldsymbol{v}$, we define the (scalar-valued) jump $[\![\boldsymbol{v}]\!]$ on each facet $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$ as
@@
  • any facet $e$ of $E$ has diameter $h_{e}$ $\geq\,\varrho\,h_{E}$;
  • for $d=3$, every face $e$ is star-shaped with respect to a ball $B_{e}$ of radius $\geq\,\varrho\,h_{e}$;
  • the mesh is quasi-uniform, i.e., any element has diameter $h_{E}\geq\varrho h$.
We introduce some basic spaces that are useful below. Given two positive integers $n$ and $m$, and $p\in[0,+\infty]$, for any $E\in\Omega_{h}$ we define
@@
    with obvious extensions for functions defined on a facet $\Pi_{n}^{0,e}\colon L^{2}(e)\to\mathbb{P}_{n}(e)$, and for vector-valued functions $\boldsymbol{\Pi}^{0,E}_{n}\colon[L^{2}(E)]^{2}\to[\mathbb{P}_{n}(E)]^{2}$ and $\boldsymbol{\Pi}^{0,e}_{n}\colon[L^{2}(e)]^{2}\to[\mathbb{P}_{n}(e)]^{2}$;
@@
$\|\varphi-\Pi^{0,E}_{n}\varphi\|_{W^{m,p}(E)}\lesssim h_{E}^{s-m}|\varphi|_{W^{s,p}(E)}\qquad$ $s,m\in\mathbb{N}$, $m\leq s\leq n+1$, $p=1,\dots,\infty$,
@@
$\|\nabla\varphi-\boldsymbol{\Pi}^{0,E}_{n}\nabla\varphi\|_{m,E}\lesssim h_{E}^{s-1-m}|\varphi|_{s,E}\qquad$ $s,m\in\mathbb{N}$, $m+1\leq s\leq n+2$.
We also have a counterpart of this result for projections defined on a facet.
@@
Let $e\in\mathcal{E}_{h}^{0}$ be an internal facet and let $E,E^{\prime}\in\Omega_{h}$ be such that $e\subset\partial E\cap\partial E^{\prime}$. Moreover, let $\Pi^{0,e}_{n}$ denote the $L^{2}(e)$-orthogonal projection onto $\mathbb{P}_{n}(e)$. Then, for every sufficiently smooth function $\varphi$ defined on $E\cup E^{\prime}$, we have that
@@
Proof.
Considering the element $E$, for $m,n\in\mathbb{N}$, with $11\leq m\leq n+$ 1, we have
$\|\varphi-\Pi^{0,e}_{n}\varphi\|_{0,e}\leq\|\varphi-\Pi^{0,E}_{n}\varphi\|_{0,e}$ $\lesssim\bigl{(}h_{E}^{-1/2}\|\varphi-\Pi^{0,E}_{n}\varphi\|_{0,E}+h_{E}^{1/2}|\varphi-\Pi^{0,E}_{n}\varphi|_{1,E}\bigr{)}$
$\lesssim h_{E}^{m-\frac{1}{2}}|\varphi|_{m,E}.$
Since the same estimate applies also to the element $E^{\prime}$, bound (10) follows. ∎
@@ 2.3 Virtual Element Spaces
Given an element $E\in\Omega_{h}$ with $n_{E}$ facets, and a positive integer $k$, we define the space
@@
  • $\mathcal{E}^{E,k}$: the moments up to the order $k-1$ on each facet $e\subset\partial E$:
    $\mu_{E,e}^{{\boldsymbol{\ell}}}(v_{h})\coloneqq\dfrac{1}{|e|}\int_{e}v_{h}\left(\dfrac{{\bf s}-{\bf s}_{e}}{h_{e}}\right)^{{\boldsymbol{\ell}}}{\rm d}s\,,\quad|{\boldsymbol{\ell}}|\leq k-1,$ (12)
    where ${\bf s}$ is expressed in the local $d-1$ coordinates on $e$, ${\bf s}_{e}$ is the barycenter of $e$, and ${\boldsymbol{\ell}}\in\mathbb{N}^{d-1}$ is a multi-index with $d-1$ components;
@@
    where $\boldsymbol{x}_{E}$ is the barycenter of $E$, and $\boldsymbol{\alpha}\in\mathbb{N}^{d}$ is a multi-index with $d$ components.
In particular, for $k=1$, the basis functions for the space $V_{h}^{k,nc}(E)$ are associated only with the facet DoFs; for $k\geq 2$, there are also the interior DoFs. In Figure 1, we show the DoFs on a penthagon for $k=1,2,3$.
@@
$N_{E}=\begin{cases}k\,n_{E}+\dfrac{(k-1)(k-2)}{2}\qquad&\text{if $d=2$}\,,\\[8.5359pt] \dfrac{(k+1)k}{2}n_{E}+\dfrac{(k+1)k(k-1)}{6}\qquad&\text{if $d=3$}\,.\end{cases}$
@@
$\int_{e}[\![v_{h}]\!]\cdot\boldsymbol{n}^{e}\,q\,{\rm d}s=0\quad\forall q\in\mathbb{P}_{k-1}(e)\quad\forall e\in{\mathcal{E}}_{h}^{0}\bigr{\}}\,.$
We remark that we do not impose full continuity across the element boundaries. On each interior facet, we only require that the moments up to order $k-1$ are preserved. Therefore, $V_{h}^{k,nc}(\Omega_{h})\not\subset H^{1}(\Omega)$. The global DoFs are as follows:
@@
    $\mu_{e}^{\boldsymbol{\ell}}(v_{h})\coloneqq\dfrac{1}{|e|}\int_{e}v_{h}\left(\dfrac{{\bf s}-{\bf s}_{e}}{h_{e}}\right)^{{\boldsymbol{\ell}}}{\rm d}s\,,\quad|{\boldsymbol{\ell}}|\leq k-1,$
    where ${\boldsymbol{\ell}}\in\mathbb{N}^{d-1}$ is a multi-index with $d-1$ components;
@@
    where $\boldsymbol{\alpha}\in\mathbb{N}^{d}$ is a multi-index with $d$ components.
@@
$N=\begin{cases}k\,|\mathcal{E}|+\dfrac{(k-1)(k-2)}{2}\,|\Omega_{h}|\qquad&\text{if $d=2$}\,,\\[8.5359pt] \dfrac{(k+1)k}{2}\,|\mathcal{E}|+\dfrac{(k+1)k(k-1)}{6}\,|\Omega_{h}|\qquad&\text{if $d=3$}\,,\end{cases}$
where $|\mathcal{E}|$ and $|\Omega_{h}|$ are the number of facets and of elements, respectively. Letting $v_{\mathcal{I}}$ be the DoFs-interpolant of a sufficiently smooth function $v$ into $V_{h}^{k,nc}(\Omega_{h})$, we recall the following interpolation error estimate, see [1, 17].
@@
Under assumption (A1), for any $v\in H^{1}(\Omega)\cap H^{s+1}(\Omega_{h})$, there exists $v_{\mathcal{I}}\in V_{h}(\Omega_{h})$, such that for all $E\in\Omega_{h}$,
@@ 2.4 Virtual Element Forms and the Discrete Problem
The local bilinear form $a^{E}(\cdot,\cdot)$ is discretized by replacing the gradient with its projection into the space $[\mathbb{P}_{k-1}(E)]^{d}$. We define the bilinear form $a_{h}^{E}(\cdot,\cdot)\colon V_{h}^{k,nc}(E)\times V_{h}^{k,nc}(E)\to\mathbb{R}$ as follows:
$\begin{split}a_{h}^{E}(u_{h},v_{h})&:=\int_{E}{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{h}\cdot{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\,{\rm d}E+\mathcal{S}^{E}\bigl{(}(I-{\Pi^{\nabla,E}_{k}})u_{h},\,(I-{\Pi^{\nabla,E}_{k}})v_{h}\bigr{)}\,,\end{split}$ (16)
@@
The reasons behind this bilinear form will become clear in Lemma 3.3 below and are related to the nonconformity of the method. Following [16, 15], we also introduce a VEM version of the local CIP-stabilization form, in order to ensure stability in the advection-dominated regime. It is defined as
@@
where $\kappa_{e}$ and $\kappa_{E}$ are positive constants to be chosen. We stress that, whenever $\boldsymbol{\beta}={\bf 0}$ inside an element $E$, the corresponding parameters $\gamma_{e}$ and $\gamma_{E}$ are zero, and hence $J_{h}^{E}(u_{h},v_{h})=0$. In order to complete the definition of our method, we need to impose the boundary conditions. We do this using a Nitsche-type technique, which locally consists of adding the form
@@
see [26, 22] for more details. By adding these terms together, we define the complete local bilinear form $\mathcal{A}_{\rm cip}^{E}\colon V_{h}^{k,nc}(E)\times V_{h}^{k,nc}(E)\to\mathbb{R}$ as
$\mathcal{A}_{\rm cip}^{E}(u_{h},v_{h})=\varepsilon a_{h}^{E}(u_{h},v_{h})+b^{E}_{h}(u_{h},v_{h})+\sigma c_{h}^{E}(u_{h},v_{h})+d^{E}_{h}(u_{h},v_{h})+\mathcal{N}_{h}^{E}(u_{h},v_{h})+J_{h}^{E}(u_{h},v_{h})\,.$ (24)
By summing over all the mesh elements, we obtain the global versions of the bilinear forms
@@ 3 Theoretical Analysis
Due to the nonconformity of the virtual element space and due to the projections entering formulation (27), the solution $u$ of the continuous problem (2) does not solve the discrete problem (27). However, if $u\in H^{2}(\Omega)\cap H_{g}^{1}(\Omega)$, it solves the following problem, which is strictly connected to (27):
@@ 3.1 Preliminary results
Since $\Delta v_{h}$ is polynomial we have (see for instance [6] [^1] Several of the results we quote have been proven in the references for $d=2$. Nevertheless, these results can be extended to $d=3$ in a straightforward manner.)
@@
$\pi v=\sum_{e\in\mathcal{E}_{h}}\sum_{|{\boldsymbol{\ell}|\leq k-1}}\mu_{e}^{{\boldsymbol{\ell}}}(v)\varphi_{e}^{{\boldsymbol{\ell}}}+\sum_{E\in\Omega_{h}}\sum_{|\boldsymbol{\alpha}|\leq k-2}\mu^{\boldsymbol{\alpha}}_{E}(v)\varphi_{E}^{\boldsymbol{\alpha}}\,,$ (45)
where $\{\varphi_{e}^{{\boldsymbol{\ell}}}\}$ is the set of basis functions associated to the DoFs at the mesh skeleton $\mathcal{E}^{k}$, and $\{\varphi_{E}^{\boldsymbol{\alpha}}\}$ is the set of basis functions associated to the interior DoFs. For an interior facet $e\in\mathcal{E}_{h}^{0}$, $e\subset\partial E\cap\partial E^{\prime}$, the coefficient $\mu_{e}^{{\boldsymbol{\ell}}}(v)$ is defined as
$\mu_{e}^{{\boldsymbol{\ell}}}(v)\coloneqq\dfrac{1}{2}\left(\mu_{E,e}^{{\boldsymbol{\ell}}}(v)+\mu_{E^{\prime},e}^{{\boldsymbol{\ell}}}(v)\right),$ (46)